### v1

@@ 4.3.2 Merging
From the left panel of Fig. 5 we realized that the model is subject to the problem of mode collapse, where it just tends to predict values close to the mean of the parent distribution. To address the issue, we merged data from the MDPL2 and Virgo $\rm\Lambda CDM$ ($\rm z=0$) simulations. Despite slight differences in simulation parameters, the turnaround radius should theoretically be ascertainable regardless of these minor variances. The outcomes of this merged data analysis are presented in the middle panel of Fig. 5. The achieved $\rm R^{2}$ Score of $0.57$, coupled with the graphical comparison of predicted versus true values, strongly suggests a significant correlation.
For additional context, we have analyzed various stacked images by categorizing the $\rm R_{ta}$ values into nine distinct bins. These are illustrated in Fig. 8. It is important to highlight that identifying the $\rm R_{ta}$ values based solely on these images poses a challenge. This difficulty is especially pronounced for values that fall outside the $\rm R_{ta}$ range of $4.3-10.6\;\rm Mpc$.
@@ 4.3.3 Interpretability and information inside $\rm R_{200}$
To better understand our findings, we analyzed the Shapley values of the pixels in a selection of test images, as depicted in Fig. 6. Our observations reveal a notable trend: a higher standard deviation in velocity in areas around the center tends to lead the algorithm to predict a lower $\rm R_{ta}$ value, and conversely, a lower velocity standard deviation suggests a higher predicted $\rm R_{ta}$. This pattern appears to support our initial hypothesis of a correlation between velocity dispersion and the mass of the central halo. However, it is crucial to note that these conclusions are drawn from just two specific examples. While Shapley values offer valuable insights, they are indicative rather than definitive. Nevertheless, a common pattern observed across most of the analyzed test images is the algorithm’s apparent emphasis on the central region of the cluster in making its predictions.
Refer to captionRefer to caption
Figure 6: Shapley values of 2 random test projections (upper and lower panel) from the merged and stacked MDPL2 and Virgo $\Lambda$CDM data. Left panel: Standard deviation of the halo velocity in each pixel. Right panel: Shapley values in each pixel (pink: higher prediction; blue: lower prediction). The red circle represents the $\rm R_{200}$, while the black and purple circles represent the true and the predicted $\rm R_{ta}$ respectively. Lower dispersion is always prominent in the central area due to the virialization of the matter surrounding the central perturbation. However, the extent of this lower dispersion area seems to indicate to the model where the turnaround radius should be. In particular, in the first image, the structure seems to have higher dispersion in areas far from the center, compared to the second one which has lower values. The high dispersion seems to indicate to the model that the turnaround radius should be lower while the lower dispersion indicates the opposite.
@@ 5 Conclusions
The turnaround density of galaxy clusters has been established as a novel cosmological probe and a means to further test the current cosmological model (Pavlidou & Tomaras 2014; Pavlidou et al. 2020; Korkidis et al. 2020, 2023). In this work, we explore a method to estimate the turnaround radius of galaxy clusters based on their projected distributions on the plane of the sky. Using cosmological N-body simulations, we generate idealized observational data—incorporating mass and velocity information without uncertainties—within two “slices” of varying thicknesses, corresponding to what is achievable in spectroscopic and photometric surveys. The “true” turnaround radii are derived from the velocity profiles surrounding the central overdensities of galaxy clusters in the simulations. Then, a Convolutional Neural Network (CNN) is trained to predict the turnaround radius, utilizing various combinations of training and test data from different simulations, redshifts, and cosmologies.
@@
When the model is trained on individual clusters we find that the mass is the most important information we have for the prediction of the turnaround radius. This at first may appear counter-intuitive as the turnaround is defined and computed by the velocity profiles of the dark matter particles around the central halo. However, it is the mass of the cluster that contributes to the gravitational dynamics defining the turnaround radius. Furthermore, we find that the mass of the central, more massive halo, is almost entirely contributing to this prediction. In the context of the findings of Korkidis & Pavlidou (2024), this is not surprising: $\rm R_{ta}$ is an overdensity radius, so its value is entirely determined at a given redshift and for a given cosmology by the enclosed mass; and that enclosed mass has been shown analytically to scale with the central, collapsed (virial) mass of the cluster.
@@
The velocity dispersion of the galaxies around the cluster can also be used as an information carrier, since it correlates with the enclosed mass. However, the small number of halos which are large enough to be observed within some “neighboring” region around a single cluster is not adequate for an accurate calculation of the velocity dispersion. Here, we counteracted this effect by using “stacked” images combining several projections. The velocity dispersion of the halos around a cluster holds important promise for analyses of this type, and it is worth exploring further in future work.
@@
As expected, contamination from distant halos that results from the use of photometric redshifts is affecting the results. Spectroscopic observations would generally result to a more accurate prediction of the turnaround radius.
@@
All of the models, when tested on data that were acquired from different redshifts and different cosmologies were not successful in their predictions. For the models trained with the mass information this is expected since the turnaround radius scales with the mass for a specific redshift and cosmology. For the velocity trained models the same pattern appears, and our attempts for a straightforward rescaling that would allow the model to generalize were not successful, although they did improve performance somewhat (see appendix E.2), so this approach with different rescaling techniques might in fact hold the most promise for a machine-learning approach to identify the turnaround scale in a cosmology-independent way.
@@
Future investigations could benefit from larger and higher-resolution datasets for training, with or without stacking. A proper Bayesian treatment or bias-reduction techniques such as the one described in Lin et al. (2022) could become particularly useful. Furthermore, more complex models e.g. with the use of inception layers in the architecture of the neural network, possibly combined with another kernel choice in smoothing the density field, or the use of Graph Neural Networks (GNNs) might be able to reveal additional information encoded in the mass and velocity information of galaxy clusters.
@@ Acknowledgements.
The CosmoSim database used in this paper is a service by the Leibniz-Institute for Astrophysics Potsdam (AIP). The MultiDark database was developed in cooperation with the Spanish MultiDark Consolider Project CSD2009-00064. The authors gratefully acknowledge the Gauss Centre for Supercomputing e.V. (www.gauss-centre.eu) and the Partnership for Advanced Supercomputing in Europe (PRACE, www.prace-ri.eu) for funding the MultiDark simulation project by providing computing time on the GCS Supercomputer SuperMUC at Leibniz Supercomputing Centre (LRZ, www.lrz.de). The Bolshoi simulations have been performed within the Bolshoi project of the University of California High-Performance AstroComputing Center (UC-HiPACC) and were run at the NASA Ames Research Center. The rest of the simulations in this paper were carried out by the Virgo Supercomputing Consortium using computers based at Computing Centre of the Max-Planck Society in Garching and at the Edinburgh Parallel Computing Centre. The data are publicly available at www.mpa-garching.mpg.de/galform/virgo/int_sims
Appendix A Additional data values distributions
@@
Refer to captionRefer to caption
Figure 7: Histogram of $\rm R_{ta}$ values from MDPL2 form $z=0.49$ and $z=1.03$ (left panel) and Virgo from OCDM and $\rm\Lambda CDM$ cosmology (right panel), similarly to Fig.1. It is apparent that the distribution moves as we change the redshift or $\rm\Omega_{m}$ and $\rm\Omega_{\Lambda}$, which is expected due to the dependence of $\rm\rho_{ta}$ on the cosmological setting discussed throughout the paper and proved in Pavlidou et al. (2020).
@@ Appendix B Architecture of the Neural Network
The NN architecture used can be seen in Table. 2. During training, the Mean Absolute Error was used to evaluate the model’s performance on the validation set. In addition to this basic architecture, the hyperparameters were tuned in each case to obtain the best result. In most of the cases the following additional elements were used: Activation function: ReLU, Padding: same, Kernel initializer: He initialization, Learning rate: $10^{-3}$, optimizer: Adam (Kingma & Ba 2014). In some of the cases l1 and l2 regularizations were also implemented.
Table 2: Characteristics of each layer of the CNN architecture. Columns are: name of the layer, input layer, size of the convolution kernel (in pixels), size (height $\times$ width in pixels) and number of the resulting feature maps.
Layer Inputs Kernel size $h\times w$feature maps
@@
Appendix C Velocity dispersion
@@
In Fig. 8 we demonstrate that acquiring $\rm R_{ta}$ from stacked velocity profiles is not straightforward. When computing the velocity dispersion images based on velocity data from 9 $\rm R_{ta}$ bins, the location of the turnaround scale is not identifiable “by eye” in the images.
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
@@
Figure 8: Stacked velocity dispersion images of the Merged MDPL2 and Virgo $\rm\Lambda CDM$ data. The turnaround values of the dataset have been segmented into 9 bins. All of the corresponding clusters’ projections of line-of-sight velocities, that belong in the same bin category, have been stacked on top of each other to create 9 images, calculating the standard deviation in each pixel. On the top of each image is the mean value of the turnaround radii in each bin which is represented with a black circle in each image. Red circles represent the mean $\rm R_{200}$ in each case. There is no apparent way of predicting the $\rm R_{ta}$ based on these images “by eye”, especially outside of the $\rm R_{ta}$ range $4.3-10.6\;\rm Mpc$.
Appendix D Stacked Velocity dispersion based on $\rm R_{200}$
@@
Here (Fig. 9) we present the results that were computed in the same way as with Fig. 5 but this time with stacking of the images based on the $\rm R_{200}$ of the clusters. The results of the middle and right panels seem to demonstrate similar success while the left panel (only using MDPL2 data) seems to be better in this case. The reason behind this is probably due to better tracing of the global minimum during training.
Refer to captionRefer to captionRefer to caption
@@
Figure 9: Comparison between predicted and true values of the $\rm R_{ta}$ for models using the ”stacked” images from 3000 bins of $\rm R_{200}$ from: the MDPL2 data (left panel); the merged MDPL2 with Virgo $\rm\Lambda CDM$ data (middle panel); the merged MDPL2 with Virgo $\rm\Lambda CDM$ data with removed information inside the $\rm R_{200}$ of the central overdensity (right panel). $\rm R^{2}$ scores are 0.5, 0.57 and 0.63 respectively. The colors represent the number density of the plotted points, calculated using a Gaussian kernel density estimate. It is apparent that the merging of the 2 datasets significantly improves the performance of the model even without the central halos’ velocity information.
Appendix E CNNs vs “Classical” astrophysical approach
We have established that significant information about the turnaround radius is encoded in the velocity dispersion of galaxies in our mock observations, in the form the stacked images of our merged dataset (MDPL2 and Virgo $\Lambda$CDM). This was achieved by the somewhat unconventional use of 2D pixelized projections processed by CNNs. In the realm of astrophysics and cosmology, a more classical and commonly used approach would be to deduce information from the radial velocity profiles of the galaxy clusters. Here, we employ this method of more traditional compression, compare it with our previous results, and explore its limitations for generalizing to other redshifts.
E.1 Baseline model
@@
As our more traditional baseline we try to train a simple model, based on the mass and velocity radial profiles of the projections, with the XGBoost algorithm (Chen & Guestrin 2016). Here, instead of the $25\times 25$ image creation described in Sect. 3.1 (2D array), we created a new dataset by calculating the mass column density and velocity dispersion in 25 radial bins in each projection (1D array). Consequently, each training instance has 25 features with the target variable being the turnaround radius.
@@
Using the mass column density MDPL2 data, the model performs with the same success, having an $\rm R^{2}$ score of $0.58$ as seen in the left panel of Fig. 10, where we plot the predicted vs true values in the same format as before. Again, this was an expected result since the model relies almost completely on the mass of the central cluster, which is not lost by the new type of compression.
Refer to captionRefer to caption
@@
A more interesting test is to compare the baseline model with the velocity dispersion data. After training it on the merged stacked dataset (which was previously the best candidate), testing resulted in an $\rm R^{2}$ score of $0.45$, as seen in the right panel of Fig. 10. In this case, the original CNN outperforms this model ($\rm R^{2}$ score: $0.57$, middle panel of Fig. 5). The underlying reason of this behavior should be the more severe compression. This compression introduces a symmetry to our data which is not necessarily present in their original form. One important symmetry breaking component is the inclusion of structures outside of the 25 Mpc radius from the velocity cut, which the CNN seems to be able to discard. It should be noted that a merged dataset with 4 copies of each projection, each rotated by 90 degrees, was performing worse even with the use of the CNN, enhancing our previous argument.
E.2 Redshift normalization
At this point the superiority of the original CNN is already substantiated from the previous section. The only caveat is that our original model is not successful when tested on other cosmological settings. For this reason we make a final attempt to counter this issue for the velocity dispersion data. In particular, we try to normalize the velocity profiles in each redshift with the background universe, in order to let the model generalize.
Subsequently, we compute the mean value of the velocity dispersion of the last radial bin of all the projections of the training dataset in the same redshift. Then, we try 2 normalizations: dividing or subtracting the values of the bins with the mean value. We train the model with the MDPL2 and the Merged data on $z=0$ and test on $z\approx 0.5$ and $z\approx 1.0$ The resulting $\rm R^{2}$ scores can be seen in Table 3.

### v2

@@ 4.3.2 Merging
From the left panel of Fig. 5, we realized that the model was subject to the problem of mode collapse, where it just tends to predict values close to the mean of the parent distribution. To address the issue, we merged data from the MDPL2 and Virgo $\rm\Lambda CDM$ ($z=0$) simulations. Despite slight differences in simulation parameters, the turnaround radius should theoretically be ascertainable regardless of these minor variances. The outcomes of this merged data analysis are presented in the middle panel of Fig. 5. The achieved $\rm R^{2}$ score of $0.57$, coupled with the graphical comparison of predicted versus true values, strongly suggests a significant correlation.
For additional context, we analyzed various stacked images by categorizing the $\rm R_{ta}$ values into nine distinct bins. These are illustrated in Fig. 9. It is important to highlight that identifying the $\rm R_{ta}$ values based solely on these images poses a challenge. This difficulty is especially pronounced for values that fall outside the $\rm R_{ta}$ range of $4.3-10.6\;\rm Mpc$.
@@ 4.3.3 Interpretability and information inside $\rm R_{200}$
To better understand our findings, we analyzed the Shapley values of the pixels in a selection of test images, as depicted in Fig. 6. Our observations revealed a notable trend: a higher standard deviation in velocity in areas around the center tends to lead the algorithm to predict a lower $\rm R_{ta}$ value, and conversely, a lower velocity standard deviation suggests a higher predicted $\rm R_{ta}$. This pattern appears to support our initial hypothesis of a correlation between velocity dispersion and the mass of the central halo. However, it is crucial to note that these conclusions are drawn from just two specific examples. While Shapley values offer valuable insights, they are indicative rather than definitive. Nevertheless, a common pattern observed across most of the analyzed test images was the algorithm’s apparent emphasis on the central region of the cluster in making its predictions.
Figure 5: Comparison between predicted and true values of the $\rm R_{ta}$ for models using the stacked velocity-dispersion images from 3000 bins of $\rm R_{ta}$ from: the MDPL2 data (left panel); the merged MDPL2 with Virgo $\rm\Lambda CDM$ data (middle panel); the merged MDPL2 with Virgo $\rm\Lambda CDM$ data with removed information inside the $\rm R_{200}$ of the central overdensity (right panel). $\rm R^{2}$ scores and the mean absolute errors (MAEs) are shown at the bottom right of each plot. The colors represent the number density of the plotted points, calculated using a Gaussian kernel density estimate. It is apparent that the merging of the two datasets significantly improves the performance of the model even without the central halos’ velocity information.
[…]
@@ 5 Conclusions
The turnaround density of galaxy clusters has been established as a novel cosmological probe and a means to further test the current cosmological model (Pavlidou & Tomaras 2014; Pavlidou et al. 2020; Korkidis et al. 2020, 2023). In this work, we explored a method to estimate the turnaround radius of galaxy clusters based on their projected distributions on the plane of the sky. Using cosmological N-body simulations, we generated idealized observational data—incorporating mass and velocity information without uncertainties—within two “slices” of varying thicknesses, approximating observational uncertainties of spectroscopic and photometric surveys. The “true” turnaround radii were derived from the velocity profiles surrounding the central overdensities of galaxy clusters in the simulations. Then, a CNN was trained to predict the turnaround radius, utilizing various combinations of training and test data from different simulations, redshifts, and cosmologies.
@@
When the model was trained on individual clusters, we find that the mass is the most important information we have for the prediction of the turnaround radius. This at first may appear counterintuitive, as the turnaround is defined and was computed by the velocity profiles of the dark matter particles around the central halo. However, it is the mass of the cluster that contributes to the gravitational dynamics defining the turnaround radius. Furthermore, we find that the mass of the central, more massive halo was almost entirely contributing to this prediction. In the context of the findings of Korkidis & Pavlidou (2024), this is not surprising: $\rm R_{ta}$ is an overdensity radius, so its value is entirely determined at a given redshift and for a given cosmology by the enclosed mass; and that enclosed mass has been shown analytically to scale with the central, collapsed (virial) mass of the cluster.
@@
The velocity dispersion of the galaxies around the cluster can also be used as an information carrier, since it correlates with the enclosed mass. However, the small number of halos that were large enough to be observed within some “neighboring” region around a single cluster was not adequate for an accurate calculation of the velocity dispersion. We counteracted this effect by using stacked images combining several projections. The velocity dispersion of the halos around a cluster holds important promise for analyses of this type, and it is worth exploring further in future work, giving a clear direction to the research community for predicting $\rm R_{ta}$.
@@
As expected, contamination from distant halos that results from the use of thicker slices affects the results. Thinner slices, corresponding to spectroscopic observations, would generally result in a more accurate prediction of the turnaround radius.
@@
All of the models, when tested on data that were acquired from different redshifts and different cosmologies, were not successful in their predictions. For the models trained with the mass information, this is expected since the turnaround radius scales with the mass for a specific redshift and cosmology. For the velocity trained models the same pattern appeared, and our attempts for a straightforward rescaling that would allow the model to generalize were not successful, although they did improve performance somewhat (see appendix F.2), so this approach with different rescaling techniques might in fact hold the most promise for a machine-learning approach to identify the turnaround scale in a cosmology-independent way.
Nonetheless, out-of-distribution predictions are a well-known challenge in simulations, particularly in machine learning based approaches. In practice, all models and simulations contain inaccuracies, yet it is assumed they approximate reality, up to some extent. There are works trying to address this issue, for example, in Jo et al. (2025) they examine the possibility of generalizing, using different cosmological simulators with different subgrid physics, for the specific task of inferring $\Omega_{m}$ and $\sigma_{8}$ from HI density slices, using a de-classifier. This is a complicated task that requires a careful examination of the latent spaces involved, using dimensionality reduction techniques, and with many caveats that are beyond the scope of this work.
[…]
@@
Future investigations could benefit from larger and higher-resolution datasets for training, with or without stacking, addressing or ignoring the problem of model misspecification. A proper Bayesian treatment or bias-reduction techniques such as the one described in Lin et al. (2022) could become particularly useful. Furthermore, more complex models—for example, with the use of inception layers in the architecture of the neural network—possibly combined with another kernel choice in smoothing the density field or further exploration of GNNs, might be able to reveal additional information encoded in the mass and velocity information of galaxy clusters.
@@ Acknowledgements.
The CosmoSim database used in this paper is a service by the Leibniz-Institute for Astrophysics Potsdam (AIP). The MultiDark database was developed in cooperation with the Spanish MultiDark Consolider Project CSD2009-00064. The authors gratefully acknowledge the Gauss Centre for Supercomputing e.V. (www.gauss-centre.eu) and the Partnership for Advanced Supercomputing in Europe (PRACE, www.prace-ri.eu) for funding the MultiDark simulation project by providing computing time on the GCS Supercomputer SuperMUC at Leibniz Supercomputing Centre (LRZ, www.lrz.de). The Bolshoi simulations have been performed within the Bolshoi project of the University of California High-Performance AstroComputing Centre (UC-HiPACC) and were run at the NASA Ames Research Centre. The rest of the simulations in this paper were carried out by the Virgo Supercomputing Consortium using computers based at Computing Centre of the Max-Planck Society in Garching and at the Edinburgh Parallel Computing Centre. The data are publicly available at www.mpa-garching.mpg.de/galform/virgo/int_sims
Appendix A Additional data distributions
@@
Figure 7: Histogram of $\rm R_{ta}$ values from MDPL2 form $z=0.49$ and $z=1.03$ (left panel) and Virgo from OCDM and $\rm\Lambda CDM$ cosmology (right panel), similarly to Fig.1. The distribution moves as we change the redshift or $\rm\Omega_{m}$ and $\rm\Omega_{\Lambda}$, which is expected due to the dependence of $\rm\rho_{ta}$ on the cosmological setting discussed throughout the paper and proved in Pavlidou et al. (2020).
@@ Appendix B Architecture of the neural network
The NN architecture used can be seen in Table. 2 (implemented using Tensorflow, Abadi et al. 2015). During training, the mean absolute error was used to evaluate the model’s performance on the validation set. In addition to this basic architecture, the hyperparameters were tuned in each case to obtain the best result. In most cases, the following additional elements were used: Activation function: ReLU, Padding: same, Kernel initializer: He initialization, Learning rate: $10^{-3}$, optimizer: Adam (Kingma & Ba 2014). In some of the cases, L1 and L2 regularizations were also implemented.
Table 2: Characteristics of each layer of the CNN architecture. [^6] Columns are: name of the layer, input layer, size of the convolution kernel (in pixels), size (height $\times$ width in pixels), and number of the resulting feature maps.
Layer Inputs Kernel size $h\times w$ #feature maps
@@
Appendix C Sensitivity to image resolution and smoothing
Using the MDPL2 data (“mass,” “vel,” and “num”), we investigated the impact of image resolution and smoothing on the CNN’s performance. To assess the effect of resolution, we tested $48$ variations of the main NN architecture of Appendix B with $\rm{number\;of\;feature\;maps}\in\{16,32,64\}$, $(h,w)\in\{(2,2),(3,3),(5,5),(10,10)\}$ for both of the convolutional layers in the network and $\rm pooling\;kernel\;sizes\in\{(1,1),(2,2),(3,3),(5,5)\}$. For each resolution, the best $\rm R^{2}$ score achieved across these variations is shown in the left panel of Fig. 8. The performance remains relatively stable across resolutions, and a resolution of $25\times 25$ pixels was selected as an optimal, balancing numerical stability and information content. To evaluate the impact of smoothing, we repeated the analysis using images smoothed with a Gaussian kernel density estimate, with bandwidths set either by Silverman’s rule of thumb (Silverman 1986) or by the pixel size (right panel of Fig. 8). Smoothing did not improve the performance of the CNN. This was expected since (a) in an ideal scenario, the CNN would be able to learn its own smoothing, if beneficial, and (b) the intrinsic Poissonian nature of the data is lost, reducing information about the number of collapsed structures.
Figure 8: Resolution sensitivity test results for the base CNN model and its variations. Left panel: $\rm R^{2}$ scores of the best-performing model for each tested resolution ($\sqrt{\text{Total pixels}}\in{3,5,11,25,65,125}$), with example “mass” images shown below. Right panel: Same test applied to images smoothed with a Gaussian kernel density estimate, using either Silverman’s rule of thumb (Silverman 1986) or a bandwidth equal to the pixel size. Example “mass” images for the former are shown below. The chosen resolution (dotted purple line and shaded purple region) provides an optimal balance between predictive accuracy and numerical stability.
Appendix D Velocity dispersion
@@
In Fig. 9 we demonstrate that acquiring $\rm R_{ta}$ from stacked velocity profiles is not straightforward. When computing the velocity dispersion images based on velocity data from nine $\rm R_{ta}$ bins, the location of the turnaround scale is not identifiable “by eye” in the images.
@@
Figure 9: Stacked velocity dispersion images of the merged MDPL2 and Virgo $\rm\Lambda CDM$ data. The turnaround values of the dataset were segmented into nine bins. All of the corresponding clusters’ projections of line-of-sight velocities, which belong in the same bin category, were stacked on top of each other to create nine images, calculating the standard deviation in each pixel. On top of each image is the mean value of the turnaround radii in each bin, which is represented with a black circle in each image. Red circles represent the mean $\rm R_{200}$ in each case. There is no apparent way of predicting the $\rm R_{ta}$ based on these images “by eye,” especially outside of the $\rm R_{ta}$ range $4.3-10.6\;\rm Mpc$.
Appendix E Stacked velocity dispersion based on $\rm R_{200}$
@@
Here (Fig. 10) we present the results that were computed in the same way as with Fig. 5, but this time with stacking of the images based on the $\rm R_{200}$ of the clusters. The results of the middle and right panels seem to demonstrate similar success, while the left panel (only using MDPL2 data) seems to be better in this case. The reason behind this is probably due to better tracing of the global minimum during training.
@@
Figure 10: Comparison between predicted and true values of the $\rm R_{ta}$ for models using the stacked images from 3000 bins of $\rm R_{200}$ from: the MDPL2 data (left panel); the merged MDPL2 with Virgo $\rm\Lambda CDM$ data (middle panel); the merged MDPL2 with Virgo $\rm\Lambda CDM$ data with removed information inside the $\rm R_{200}$ of the central overdensity (right panel). $\rm R^{2}$ scores and the mean absolute errors (MAEs) are shown at the bottom right of each plot. The colors represent the number density of the plotted points, calculated using a Gaussian kernel density estimate. It is apparent that the merging of the two datasets significantly improves the performance of the model even without the central halos’ velocity information.
Appendix F CNNs versus “classical” astrophysical approach
[…]
F.1 Baseline model
@@
As our more traditional baseline, we tried to train a simple model, based on the mass and velocity radial profiles of the projections, with the XGBoost algorithm (Chen & Guestrin 2016). Here, instead of the $25\times 25$ image creation described in Sect. 3.1 (2D array), we created a new dataset by calculating the mass column density and velocity dispersion in 25 radial bins in each projection (1D array). Consequently, each training instance had 25 features with the target variable being the turnaround radius.
@@
Using the mass column density MDPL2 data, the model performed with the same success, having an $\rm R^{2}$ score of $0.58$ as seen in the left panel of Fig. 11, where we plot the predicted versus true values in the same format as before. Again, this was an expected result since the model relied almost completely on the mass of the central cluster, which is not lost by the new type of compression.
@@
A more interesting test is to compare the baseline model with the velocity dispersion data. After training it on the merged stacked dataset (which was previously the best candidate), testing resulted in an $\rm R^{2}$ score of $0.45$, as seen in the right panel of Fig. 11. In this case, the original CNN outperformed this model ($\rm R^{2}$ score: $0.57$, middle panel of Fig. 5). The underlying reason for this behavior should be the more severe compression. This compression introduces a symmetry to our data that was not necessarily present in their original form. One important symmetry-breaking component is the inclusion of structures outside of the 25 Mpc radius from the velocity cut, which the CNN seems to be able to discard. It should be noted that a merged dataset with four copies of each projection, each rotated by 90 degrees, was performing worse, even with the use of the CNN, enhancing our previous argument.
F.2 Redshift normalization
At this point, the superiority of the original CNN is already substantiated from the previous section. The only caveat was that our original model was not successful when tested on other cosmological settings. For this reason, we made a final attempt to counter this issue for the velocity dispersion data. In particular, we tried to normalize the velocity profiles in each redshift with the background Universe, in order to let the model generalize.
Subsequently, we computed the mean value of the velocity dispersion of the last radial bin of all the projections of the training dataset in the same redshift. Then, we tried two normalizations: dividing by or subtracting the values of the bins from the mean value. We trained the model with the MDPL2 and the merged data on $z=0$ and tested on $z\approx 0.5$ and $z\approx 1.0$. The resulting $\rm R^{2}$ scores can be seen in Table 3.